# Cross Validation in Stochastic Analytic Continuation

### v1

@@ IV.2 Double Guassian Spectrum
Figure 3: For the spectral function shown in Fig. 4, the average of $K=40$ cross validation annealing runs using the synthetic data with a noise level of $\sigma=10^{-5}$. Panels, symbols, and set up are the same as in Fig. 1.
@@
Figure 4: Deviation between the extra spectrum and those sampled at three different temperatures: the temperature where $a=0.5$ (blue), the temperature where $\chi^{2}_{\mathrm{val}}$ is minimized (red), and at a slightly lower temperature, corresponding to one standard deviations below this minimum (green). We note that the magnitude of the deviation is both a reflection of the error in the amplitudes and the locations of the two Gaussian peaks. Inset: the spectral function themselves. Just as in Fig. 2, the red and blue spectra are nearly indistinguishable.
@@
We preform the same test as in Sec. IV.1, and present the results in Fig. 3. In this case, both the validation minimum and the temperature range corresponding to the criterion with $a$ ranging from 0.25 and 1 are much broader, but still align quite well with one another. The higher level of uncertainty is a reflection of the inability of the unconstrained sampling scheme to reproduce this spectrum nearly as well as in the previous case, which can be seen clearly in Fig. 4. Here we again plot the spectra at the validation $\chi^{2}$ minimum (red) and at the temperature where $a=0.5$ (blue), as well as at a slightly lower temperature, corresponding to one standard deviation away from the validation minimum (green). It is interesting to note that, at the lower sampling temperature, SAC is able to reproduce certain features of the exact spectrum better than at the temperatures corresponding to the validation minimum or $a=0.5$ (such as the position of the first peak and area where the two peaks intersect), while reproducing others not as well (such as the position of the second peak). This is consistent with the broadness of the validation $\chi^{2}$ minimum, as it is not necessarily clear which is the better choice of spectra.
@@
We emphasize that using cross validation to determine the optimal sampling temperature is not necessary when running SAC with real QMC data. Rather, these test-cases show that the criterion in Eq. (11) indeed samples safely above the regime of over-fitting, while still keeping the $\chi^{2}$ at an acceptable level, and should be used as the method of fixing $\Theta$ in real use-cases. In the following sections we will explore a more practical and decisive use for our cross validation procedure.
V Optimal Spectral Parameterization
V.1 Heisenberg Chain
@@
Our second application of cross validation will be to help determine which SAC sampling parameterization to use when the exact spectral features are unknown. In Eq. (18), we model the spectrum as unconstrained $\delta$-functions with variable amplitudes and positions. However, if certain restrictions are imposed on the locations and amplitudes of the $\delta$-function, the SAC method is able to reproduce sharp spectral features present only at low temperature, such as narrow quasi-particle peaks and power-law edge singularities [10, 21, 14]. We call this class of SAC sampling schemes “constrained.”
@@
A spectrum containing a dominate $\delta$-peak followed by a smooth continuum can be reproduced using the parameterization
@@
where $A_{0}+\sum_{i}A_{i}=1$ and the constraint $\omega_{i}>\omega_{0}$ for all $i$ is imposed during the sampling process. With this parameterization, the location of the peak, $\omega_{0}$, is sampled within the program, along with all of the other $\delta$-functions, but the weight of the dominant $\delta$ peak, $A_{0}$, is fixed and optimized using a simple scan to find a $\chi^{2}$ minimum, as detailed in Ref. [14]. This type of spectral function is appropriate for describing the spectrum of the $S=1/2$ AFM Heisenberg model in two-dimensions (2D) [22], where linear spin-wave theory predicts the presence of a dominant, “single-magnon” peak [23], while an incoherent continuum at higher energy can be attributed to either multimagnon excitations or partially deconfined spinon excitations, both of which cannot be captured by convention spin-wave theory [21]. The parameterization of SAC in Eq. (19) has been used to produce results for the spectral function of the aforementioned square-lattice Heisenberg model that agreed very well with results from inelastic neutron-scattering experiments on a material considered the best physical realization of this model [21].
Another example of a sharp spectral feature that can only be resolved using a constrained SAC sampling parameterization is an edge singularity. In quantum many-body systems with fractionalized excitations, spectral weight may be spread over a range of energies, often with a power-law distribution that diverges at some frequency $\omega_{q}$, where $q$ is the total momentum of the fractionalized quasiparticle [14]. One of the most well know examples of this is the spectral function of the operator $O=S_{q}^{z}(\omega)$ in the $S=1/2$ AFM Heisenberg chain. Using the Bethe ansatz (BA) solution, it was shown that in the thermodynamic limit, this spectral function diverges as $(\omega-\omega_{q})^{-1/2}$, with a logarithmic correction [24, 25, 26].
A spectrum with this type of feature can be realized in SAC by restricting the locations of the sampled $\delta$-functions such that the distance between adjacent $\delta$’s monotonically increases [10, 14]:
@@
with the constraint $\omega_{i+1}-\omega_{i}>\omega_{i}-\omega_{i-1}$, i.e. monotonically increasing spacing. Due to entropic pressure, this parameterization naturally produces a spectrum where the mean density of the $\delta$-functions forms precisely an edge singularity with the asymptotic behavior $(\omega-\omega_{q})^{-1/2}$. Away from this edge, the data instead dictates the exact shape of the spectrum. It also possible to resolve an edge diverging with any power $p$, that can be optimized with a scan, by properly adjusting the amplitudes $A_{i}$, as detailed in Ref. [14], but in these tests we just consider the $p=1/2$ case.
@@
In most cases, the exact features of the true spectral function are not known, and one must test out these various parameterizations. However, relying on the value of $\chi^{2}$ alone to compare the resulting spectra is not always reliable, since the SAC method can produce spectra with an acceptable $\chi^{2}$ values even when the improper parameterization is being used (a consequence of the “ill-posed” problem). Our cross validation procedure can be used as an unbiased method to determine which parameterization is best amongst the group of parameterizations considered, i.e. for “model selection” [17].
We demonstrate the ability to discriminate between models using real QMC data, generated using the stochastic series expansion (SSE) method [27], for the $S=1/2$ AFM Heisenberg chain. We follow the procedure outlined in Sec. III, again with $K=20$ validation sets with an error level of $\sigma\sim 10^{-5}$. The system size used here is $L=512$, and the QMC simulation was run using the inverse temperature $\beta=1024$, which is large enough to access ground state properties. While the predicted power-law form holds for all momenta $q$ in $S_{q}^{z}(\omega)$, we consider just $q=\pi/2$, as a test case.
@@
Figure 5: Cross validation results for $S(q=\pi/2,\omega)$ of the $S=1/2$ Heisenberg chain ($L=512$, $\beta=1024$). The sampling (a) and validation (b) $\chi^{2}$ values, normalized by the number of $\tau$ points, are averaged over all rotations of the sampling and validation data sets (twenty in total). In panel (c), we plot the validation $\chi^{2}$ as a function of the sampling $\chi^{2}$, to compare the three parameterizations at fixed sampling $\chi^{2}$ values.
For this test, we repeat the cross validation process, rotating which of the $K+1=21$ sets of QMC data is used for sampling, and which ones are used for validation, which we found was necessary given the quality and quantity of the QMC-generated data. The final result is the average over all of the rotations, which we use to compare each of the three parameterizations.
Fig. 5 shows the sampling (panel (a)) and validation (panel (b)) $\chi^{2}$ values versus $\Theta$ for each parameterization: unconstrained (red), single-peak (blue), and edge (green). Here, we normalize by the number of $\tau$ points (in this case 50, with $\Delta_{\tau}=0.1$) and subtract one from the normalized validation $\chi^{2}$ to account for the background contribution, $X_{2}$ in Eq. (14). Because the entropy contents of the spectra defined with each parameterizations are different, the annealing paths they each take differ, so it is also informative to compare the validation $\chi^{2}$ values at the same value of the sampling $\chi^{2}$. This is done in panel (c), where we plot the validation $\chi^{2}$ versus the sampling $\chi^{2}$ throughout the annealing runs.
@@
Figure 6: Spectra corresponding to the three parameterizations in Fig. 5, unconstrained (a), single-peak with $A_{0}=0.50$ (b), and edge (c). For each parameterization, the spectra are sampled at the temperatures corresponding to the criterion in Eq. (11), using their respective sampling $\chi^{2}$ curves. For the unconstrained sampling parameterization, we use $N\omega=2,000$ $\delta$-functions. For the single-peak parameterization we use a $N_{p}=1$ leading $\delta$-function with a relative weight of $A_{0}=0.5$ and $N_{c}=2,000$ continuum $\delta$-functions. For the edge parameterization, we use $N_{\omega}=320$ $\delta$-functions. In each panel, we also plot the BA results (black curves), which have also been subject to smoothing [28], resulting in more rounded peaks. Insets: same spectra, but on zoomed in scales.
@@
While it is difficult to glean any information about the relative performance of the three parameterizations from the sampling $\chi^{2}$, the validation $\chi^{2}$ paints a much different picture. At each temperature, the the edge parameterization has the lowest validation $\chi^{2}$, followed by the single peak case, and then unconstrained sampling. This would suggest that the edge is indeed the most appropriate parameterization, in agreement with the BA solution for this model. This conclusion can also be made when comparing the values of $\chi_{\mathrm{val}}^{2}$ at fixed $\chi^{2}$, panel (c). Here, the single peak and edge parameterizations preform equally as well up until the normalized sampling $\chi^{2}$ reaches $\sim 1.27$, at which point the validation $\chi^{2}$ curve for the edge parameterizations dips below that of the single peak. The $\chi^{2}$ values corresponding to the criterion in Eq (11) with $a=0.5$ are 1.26, 1.30 and 1.30 for the unconstrained, peak, and edge parameterization, respectively, which are the values for the spectra in Fig. 6. We note, however, that the spectra do not change significantly when sampled at slightly lower temperatures, corresponding to the minima of each $\chi_{\mathrm{val}}^{2}$ curve, just as in Sec. IV.1. We conclude that it is thus important take into consideration both the minimum value of each $\chi_{\mathrm{val}}^{2}$ curve and the path the curves take when making these comparisons.
@@
We now compare the corresponding spectra, shown in Fig. 6, to see if the hierarchy suggested by the validation $\chi^{2}$ is borne out in the results of the analytic continuation process. As reflected by the relative differences between the $\chi_{\mathrm{val}}^{2}$ curves, the single-peak and edge spectra share similar features, such as a lower bound near $\omega=1.5$, while the unconstrained sampling produces a spectra with considerable spectral weight below this edge. The BA result [28], shown in black, agrees very well with the edge spectra (with edges that nearly perfectly align), moderately well for the peak spectra (but with some glaring deviations), and very poorly with the unconstrained spectra, further supporting the results of our cross validation test. Here, it should be noted that the BA spectrum has been subject to significant broadening, and furthermore, the calculation of this spectrum only took into consideration two-spinon and four-spinon contributions. As such, it is known that a few percect of the total spectral weight is missing, which could at least partially explain the deviations in the tail in Fig. 6(c).
@@
While it is encouraging that in this test the edge parameterization beats out the unconstrained and single-peak, the difference between the validation $\chi^{2}$’s is likely too small to claim with any certainly that the spectrum truly contains this feature. Furthermore, for a chain of length $512$, the exact $T=0$ spectrum likely only contains $\sim$100 delta functions with significant weight, deduced by extrapolating from the results for smaller system sizes [29, 30]. With this in mind, its quite possible that a spectrum composed of a dominate peak plus a smooth continuation is actually the best representation for this finite system, while only in the thermodynamic limit would the edge parameterization truly be the most accurate. Preforming more cross validation tests using synthetic spectra, where these sharp features are exactly realized, will be necessary to gauge the reliability of this procedure.
V.2 Heisenberg Chain with Long-Range Interactions

### v2

@@ IV.1 Single Gaussian Peak
Figure 4: For the spectral function shown in Fig. 5, the average of $K=20$ cross validation annealing runs using synthetic data with a noise level of $\sigma=10^{-5}$. Panels, symbols, and set up are identical to those in Fig. 2.
IV.2 Two Gaussian Peaks
We now turn to an example where the unconstrained sampling parameterization in Eq. (18) cannot reproduce the artificial spectrum quite as closely. The spectrum we consider is composed of a pair of Gaussian peaks, one broader than the other, that intersect slightly. The area near the sharp dip, where the two peaks overlap, poses difficulties and would require a lower error level than one could reasonably achieve using real QMC-generated data to resolve fully [20]. As such, the criterion in Eq. (11) provides an estimate for the temperature corresponding to the best possible spectrum, given these limitations.
@@ IV.2 Two Gaussian Peaks
Figure 5: Deviation between the exact spectrum and those sampled at three different temperatures: the temperature where $a=0.5$ (blue), the temperature where $\chi^{2}_{\mathrm{val}}$ is minimized (red), at a slightly lower temperature, corresponding to one standard deviation below this minimum (green), and at a slightly higher temperature, corresponding to one standard deviation above the minimum (yellow). We note that the magnitude of the deviation is both a reflection of the error in the amplitudes and the locations of the two Gaussian peaks. Inset: the spectral function themselves. Just as in Fig. 3, the red and blue spectra are nearly indistinguishable.
@@
We preformed the same test as in Sec. IV.1, and present the results in Fig. 4. In this case, both the validation minimum and the temperature range corresponding to the criterion with $a$ ranging from 0.25 and 1 are much broader, but still align quite well with one another. The higher level of uncertainty is a reflection of the inability of the unconstrained sampling scheme to reproduce this spectrum nearly as well as in the previous case, which can be seen clearly in Fig. 5. Here we again plot the spectra at the validation $\chi^{2}$ minimum (red) and at the temperature where $a=0.5$ (blue), as well as at a slightly lower/higher temperature, corresponding to one standard deviation below/above from the validation minimum (green/yellow). It is interesting to note that, at the lower sampling temperature (green), SAC is able to reproduce certain features of the exact spectrum better than at the temperatures corresponding to the validation minimum or $a=0.5$ (such as the position of the first peak and area where the two peaks intersect), while reproducing others not as well (such as the position of the second peak). This is consistent with the broadness of the validation $\chi^{2}$ minimum, as it is not necessarily clear which is the better choice of spectra. Furthermore, all of the spectra shown here are acceptable reproductions of the underlying artificial spectrum; the most prominent features of the spectrum are captured well in all four cases and the deviations plotted in Fig. 5 reflect only small shifts in the locations and amplitudes of the two Gaussian peaks.
@@
We emphasize that using cross validation to determine the optimal sampling temperature is not necessary when running SAC with real QMC-generated data. Rather, these test-cases show that the criterion in Eq. (11) indeed samples safely above the regime of over-fitting, while still keeping $\chi^{2}$ at an acceptable level, and should be used as the method of fixing $\Theta$ in real use-cases. In the following sections we will explore a more practical and decisive use for our cross validation procedure, as a tool for selecting the most applicable SAC parameterization.
V Optimal Spectral Parameterization: Unconstrained Sampling
Our second application of cross validation will be to help determine which SAC parameterization to use when the exact spectral features are unknown, i.e. model selection [22, 23]. In Eq. (18), we model the spectrum as unconstrained $\delta$-functions with sampled positions (Fig. 1(a)). In addition to the positions of the $\delta$-functions, their amplitudes may also be variable and sampled, as depicted in Fig. 1(b). These two parameterizations, which we’ll refer to as equal and variable amplitudes (EA and VA, respectively), have different configurational entropies as well as different sampling efficiencies, depending on the exact form of the underlying spectrum. For the EA parameterization, the entropy is the conventional Shannon information entropy [11, 20], while for the VA parameterization, a generalized Rényi entropy is instead appropriate [21].
It was shown that if these entropies are used in MEM, the results will be identical to those from SAC, when using the corresponding unconstrained parameterization, with the caveats that $N_{\omega}$ must be large and the MEM entropy regulator $\alpha$ must be chosen such that the two output spectra have the same $\chi^{2}$ values (the choice of $\alpha$ when using the variable amplitude parameterization is slightly more subtle due to the specifics of the entropy form used) [20].
In many cases, such as the two considered in the previous section, the output spectra will depend very little, if at all, on whether amplitudes are sampled along with frequencies. More specifically, if $N_{\omega}$ is large enough, the spectral features are not very sharp, and the error level is low, the different entropy contents will only minimally affect the sampling efficiency and the underlying spectrum can be reproduced with high fidelity using either parameterization [20]. But this is not always the case, as we will explore in the following sections.
V.1 Two Equal Amplitude Gaussian Peaks
Figure 6: Cross validation results for the spectral function shown in Fig. 7. In panel (b), the sampling $\chi^{2}$ values, normalized by the number of $\tau$ points, is plotted versus the sampling temperature $\Theta$. In panel (c), the validation $\chi^{2}$ values, also normalized by the number of $\tau$ points, but now also averaged over all cross validations data sets, is plotted versus the sampling temperature $\Theta$. We have subtracted the background constant of one from the validation $\chi^{2}$ to account for the term $X_{2}$ in Eq. (14). In panel (a), the validation $\chi^{2}$ as a function of the sampling $\chi^{2}$ is plotted to allow for direct comparison of the two parameterizations at a fixed value of the sampling $\chi^{2}$.
In the first case we consider, we used an artificial spectrum composed of two equal amplitude Gaussian peaks centered at $\omega=2$ and $\omega=4$. While this spectrum is relatively simple, the proximity of the two peaks poses similar difficulties to those encountered in Sec. IV.2; the low amplitude region in between the two peaks cannot be resolved fully at this error level, which causes a ripple effect of distortions throughout the spectrum. The two unconstrained sampling schemes, EA and VA, are able to overcome these challenges to different extents, which can be seen clearly when preforming tests using artificial spectra. However, when preforming analytic continuation using QMC-generated data, one must instead rely on cross validation to asses the abilities of these two parameterizations.
Just as in Sec. IV, we generated $K+1=21$ sets of synthetic data with error level $\sigma=10^{-5}$. In this case, we used a slightly higher inverse temperature, $\beta=16$, and accordingly a larger $\tau$ spacing, $\Delta\tau=0.1$, when converting the artificial spectrum to $G(\tau)$. When defining the error level, we introduce a cutoff of $\tau_{\rm max}\approx 4$, chosen so that the relative error on $\bar{G}(\tau)$ does not exceed 10%. We ran the cross validation procedure using two different SAC parameterizations, the unconstrained sampling method with equal amplitude and variable, sampled amplitude $\delta$-functions, and compare their performance.
Figure 6 shows the results of this cross validation test. The normalized sampling $\chi^{2}$ during the annealing run is shown in panel (b) for both updating methods, equal (red) and variable (blue) amplitudes. Because the entropy contents of the spectra defined with each parameterization are different, the annealing paths they take differ, so it is also informative to compare the validation $\chi^{2}$ values at the same value of the sampling $\chi^{2}$. This is shown in Fig. 6(a) where we plot the validation $\chi^{2}$ versus the sampling $\chi^{2}$ throughout the annealing run.
Figure 7: Spectra corresponding to the two parameterizations shown in Fig. 6, equal amplitudes (panel (a)) and variable amplitudes (panel (b)). The exact spectrum is plotted in black, while the two colored spectra correspond to two different sampling temperatures: the validation $\chi^{2}$ minimum (red) and the criterion in Eq. (11) with $a=0.5$ (blue). In both cases, $N_{\omega}$ = 4,000 $\delta$-functions were used, which was chosen to ensure that there was no dependance of $S(\omega)$ on $N_{\omega}$.
For this spectrum, the validation $\chi^{2}$ reaches the lowest minimum value for the EA parameterization, which can be clearly seen in Fig. 6 panel (c), and even more clearly in panel (a). We can see why this is the case by comparing the spectra produced using both parameterizations, shown in Fig. 7. Here, we plot the spectra at two sampling temperatures, corresponding to the validation $\chi^{2}$ minima (red, with $\chi/N_{\tau}\sim 0.56$) and the criterion in Eq. 11 with $a=0.5$ (blue, with $\chi/N_{\tau}\sim 0.54$).
The EA parameterization (panel (a)) produces a spectrum where both the shapes and heights of the two Gaussian peaks are very close to those in the exact spectrum, shown in black. On the other hand, the VA parameterization (panel (b)) produces peaks that are far too sharp with amplitudes that are unequal and too large. While both parameterizations are unable to correctly reproduce the region in between the two peaks, the VA parameterization preforms far worse. The excessively tall peaks produced by the variable amplitudes parameterization may be compensation for the missing weight in this region.
Each SAC parameterization comes with its own set of entropic pressures associated with the degrees of freedom that are being sampled. In the case of the VA parameterization, it is known that these pressures overly favor sharp peaks [20]. In this example, that comes as a detriment, but in others, this can be advantageous, as we will explore in the following section.
V.2 Sharp Gaussian Peak with a Shoulder
Figure 8: Cross validation results for the spectral function shown in Fig. 9. Panels and set up are identical to those in Fig. 6
For our second test, we consider an artificial spectrum composed of a tall and sharp Gaussian peak centered at $\omega=2$, followed by a short and broad Gaussian at $\omega=3$, which forms a shoulder-like feature for the first peak. A similar artificial spectrum was used in Ref. 20 to study this exact issue, whether or not to sample amplitudes along with frequencies. It was found that including amplitude updates was necessary to resolve both the dominant peak and the shoulder for data with error levels attainable using QMC [20]. Being able to determine this using cross validation would be very valuable when the exact underlying spectrum is unknown.
Again, we generated $K+1=21$ sets of synthetic data with error level $\sigma=10^{-5}$ and use $\beta=16$ with $\Delta\tau=0.1$ (again, with $\tau_{\rm max}\approx 4$), which is the same inverse temperature used in Ref. 20 for their tests on this type of spectral function. The results of the cross validation procedure are shown in Fig. 8 and the corresponding spectral functions are shown in Fig. 9.
In this case, cross validation suggests that the VA parameterization performs better, in contrast to the previous example. Here, the validation $\chi^{2}$ reaches a lower minimum value (panel (c)), but is also lower for the majority of the annealing process, when compared at fixed values of the sampling $\chi^{2}$ (panel (a)).
Figure 9: Spectra corresponding to the two parameterizations shown in Fig. 7. Panels and set up are identical to those in Fig. 7, except an additional spectrum is plotted, corresponding to the criterion in Eq. (11) with $a=2.0$.
In Fig. 9 we plot the spectra at three different sampling temperatures for both parameterizations. The red spectra were sampled at the temperature corresponding to the validation minima, and the blue and green spectra were sampled at temperatures corresponding criterion in Eq. (11) with $a=0.5$ and $a=2.0$, respectively.
At all three sampling temperatures, the VA parameterization is able to resolve the first peak with higher fidelity. In both cases, the spectra sampled at the validation $\chi^{2}$ minima (red, with $\chi/N_{\tau}\sim 0.55$) do not reproduce the shoulder feature very well, instead displaying a ringing pattern around this structure. But as the temperature, and consequently the $\chi^{2}$ value, are increased (blue, with $\chi/N_{\tau}\sim 0.66$ and green, with $\chi/N_{\tau}\sim 1.0$), the shoulder is reproduced with much higher accuracy by the VA parameterization. This is, however, not the case for the EA parameterization, as suggested by the results shown in Fig. 8(c).
In the previous examples, there was much better agreement between the validation $\chi^{2}$ minima and the criterion in Eq. (11) than there is for this case. We explain this discrepancy by noting that at the error levels considered here, $\sigma=10^{-5}$, this spectrum is unable to be produced with nearly as high a fidelity. If the SAC results will always deviate from the exact solution in some considerable way, then the validation $\chi^{2}$ may be minimized by a spectrum that does not visually appear to match the exact spectrum quite as well as one that is produced at a slightly higher or lower sampling temperature. This is exemplified by the $a=2.0$ spectrum in Fig. 9(b); while it may visually appear to be the most accurate, the results of the cross validation procedure suggests that this is not the case. Cross validation identifies the spectrum with the corresponding $G(\tau)$ that is closest to the exact solution, and in some cases this may be a spectrum that deviates from the exact $S(\omega)$ in multiple different ways, i.e. a compensating effect that produces multiple different distortions. Even considering these limitations, it is clear that cross validation can accurately determine which unconstrained sampling scheme produces the most accurate spectrum.
VI Optimal Spectral Parameterization: Constrained Sampling
@@
There is second class of SAC parameterizations that have an even greater impact on the shape and form of the output spectra, which refer to as constrained SAC sampling schemes (Fig. 1(c)-(e)). If certain restrictions are imposed on the locations and amplitudes of the $\delta$-functions, the SAC method can reproduce sharp spectral features present only at low temperatures, such as narrow quasi-particle peaks and power-law edge singularities [16, 27, 20]. In many cases, it is known that the spectral function contains such a sharp feature, but the exact form of this feature, and thus which constrained sampling scheme to use, is unknown. In the following sections we will show how cross validation can be used to identify the most appropriate constrained SAC parameterization.
@@
A spectrum containing a dominant $\delta$-peak followed by a smooth continuum can be reproduced using the parameterization depicted in Fig. 1(b) [27]
@@
where $A_{0}+\sum_{i}A_{i}=1$ and the constraint $\omega_{i}>\omega_{0}$ for all $i$ is imposed during the sampling process. With this parameterization, the location of the peak, $\omega_{0}$, is sampled within the program, along with all of the other $\delta$-functions, but the weight of the macroscopic $\delta$-peak, $A_{0}$, is fixed and optimized using a simple scan to find a $\chi^{2}$ minimum, as detailed in Appendix A.1. This type of spectral function is appropriate for describing the spectrum of the $S=1/2$ AFM Heisenberg model in two-dimensions (2D) [28], where linear spin-wave theory predicts the presence of a macroscopic, “single-magnon” peak [29], while an incoherent continuum at higher energy can be attributed to either multimagnon excitations or partially deconfined spinon excitations (both of which cannot be captured by convention spin-wave theory [27]). The SAC parameterization in Eq. (19) has been used to produce results for the spectral function of the aforementioned square-lattice Heisenberg model that agreed very well with results from inelastic neutron-scattering experiments on a material considered the best physical realization of this model [27].
Another example of a sharp spectral feature that can only be resolved using a constrained SAC sampling parameterization is an edge singularity. In quantum many-body systems with fractionalized excitations, spectral weight may be spread over a range of energies, often with a power-law distribution that diverges at some frequency $\omega_{q}$, where $q$ is the total momentum of the fractionalized quasiparticle [20]. One of the most well know examples of this is the spectral function of the operator $O=S_{q}^{z}$ in the $S=1/2$ AFM Heisenberg chain. Using the Bethe ansatz (BA) solution, it was shown that in the thermodynamic limit, this spectral function diverges as $(\omega-\omega_{q})^{-1/2}$, with a logarithmic correction (except for at $q=\pi$, where the divergence is faster) [30, 31, 32].
A spectrum with this type of feature can be realized in SAC by restricting the locations of the sampled $\delta$-functions such that the distance between adjacent $\delta$-function’s monotonically increases (Fig. 1(c) and (d)) [16, 20]:
@@
with the constraint $\omega_{i+1}-\omega_{i}>\omega_{i}-\omega_{i-1}$, as depicted in Fig. 1(c). Due to entropic pressure, this parameterization naturally produces a spectrum where the mean amplitude density of the $\delta$-functions forms precisely an edge singularity with the asymptotic behavior $(\omega-\omega_{q})^{-1/2}$. The location of the edge, $\omega_{q}$, is not a fixed, user-inputted parameter that must be scanned over, like in the case of the leading $\delta$ weight $A_{0}$ in Eq. (19). Rather, the location of the edge adapts naturally to the data during the sampling process. Away from this edge the data dictates the exact shape of the spectrum, smoothly transitioning from the asymptotic power-law form.
It is also possible to resolve an edge diverging with any power $p$, $(\omega-\omega_{q})^{p}$ with $p<0$, by properly adjusting the amplitudes $A_{i}$ [20]. For an edge that diverge faster/slower than $S(\omega\to\omega_{q})\propto(\omega-\omega_{q})^{-1/2}$, i.e. $p$ less/greater than $-1/2$, the amplitudes of the $\delta$-functions increase/decrease as they approach the edge. This is depicted in Fig 1(d) for an edge that decays with an exponent $p=-3/4$.
@@
In most cases, one tests both the unconstrained and constrained parameterizations when preforming analytic continuation on QMC-generated data. However, relying on the value of $\chi^{2}$ alone to compare the resulting spectra is not always reliable, since the SAC method can produce spectra with acceptable $\chi^{2}$ values even when the improper parameterization is being used (a consequence of the “ill-posed” problem). Our cross validation procedure can be used as an unbiased method not only to determine whether an unconstrained or constrained sampling scheme should be used, but also to determine which of the constrained sampling scheme produces the most statistically likely spectrum.
VI.1 Heisenberg Chain
@@
Figure 10: Cross validation results for $S(q=\pi/2,\omega)$ and $S(q=3\pi/4,\omega)$ of the $S=1/2$ Heisenberg chain ($L=512$, $\beta=1024$). In panels (b) and (e), the sampling $\chi^{2}$ values, normalized by the number of $\tau$ points and averaged over all cross validations runs, is plotted versus the sampling temperature $\Theta$. In panels (c) and (f), the validation $\chi^{2}$ values, also normalized by the number of $\tau$ points (50 in this case) and averaged over all cross validations runs, is plotted versus the sampling temperature $\Theta$. We have subtracted the background constant of one from the validation $\chi^{2}$ to account for the term $X_{2}$ in Eq.(14). In panels (a) and (d), we plot the validation $\chi^{2}$ as a function of the sampling $\chi^{2}$, to allow for direct comparison of the three parameterizations at a fixed $\chi^{2}$ value.
We first demonstrate the ability to discriminate between models using real QMC data, generated using the stochastic series expansion (SSE) method [33], for the $S=1/2$ AFM Heisenberg chain. We followed the procedure outlined in Sec. III, again with $K=20$ validation sets with an error level of $\sigma=10^{-5}$. The system size used here is $L=512$, and the QMC simulation was run using the inverse temperature $\beta=1024$, which is large enough to access ground state properties. We considered $S_{q}^{z}(\omega)$ at two momenta, $q=\pi/2$ and $43\pi/$ 4, as test cases. For both $q$ values, the optimal weights of the macroscopic $\delta$-peak were determined using the method detailed in Ref. 20 and shown here in Appendix A.1 ($A_{0}=0.40$ and $0.35$ for $q=\pi/2$ and $43\pi/$ 4, respectively). Due to the logarithmic correction to the square-root divergent edge form of the spectral function for $O=S_{q}^{z}$ [30, 31, 32], the optimal power $p$ for each $q$ value considered may not necessarily be equal to $-0.50$. As we show in Appendix A.2, for $q=3\pi/4$, the optimal exponent is $p=-0.60$.
@@
Figure 11: Spectra corresponding to the three parameterizations in Fig. 10, unconstrained (a, d), single-peak with (b, e), and edge (c, f). For each parameterization, the spectra are sampled at the temperatures corresponding to the criterion in Eq. (11), using their respective sampling $\chi^{2}$ curves. For the unconstrained sampling parameterization, we use $N_{\omega}=2,000$ $\delta$-functions. For the single-peak parameterization we use the leading $\delta$-function weights $A_{0}$ shown in Fig. 10 and $N_{c}=2,000$ continuum $\delta$-functions. For the edge parameterization, we use $N_{\omega}=640$ and the exponent $p$ shown in Fig. 10. In each panel, we also plot the BA results for a chain of length 500 (black curves), which have also been subject to smoothing [34], resulting in more rounded peaks.
For this test, we repeated the cross validation process for three parameterizations (unconstrained, single-peak, and edge). We now also rotate which of the $K+1=21$ sets of QMC-generated data was used for sampling, and which ones were used for validation, and averaged the results over all of the rotations. We found that this extra step was necessary given the quality and quantity of the QMC-generated data at hand.
Fig. 10 shows the results of this test. The the normalized sampling (panels (b) and (e)) and validation (panels (c) and (f)) $\chi^{2}$ values are plotted versus $\Theta$ for each parameterization: unconstrained (red), single-peak (blue), and edge (green). Just as with the two unconstrained sampling parameterizations, the entropy contents of the spectra defined with each constrained sampling parameterization are different, so to better compare the three parameterizations, we plot the validation $\chi^{2}$ versus the sampling $\chi^{2}$ as well, shown in panels (a) and (d).
@@
While it is difficult to glean any information about the relative performance of the three parameterizations from the sampling $\chi^{2}$, the validation $\chi^{2}$ paints a much different picture. At each temperature, the the edge parameterization has the lowest validation $\chi^{2}$, followed by the single-peak case, and then unconstrained sampling. This would suggest that the edge is indeed the most appropriate parameterization, in agreement with the BA solution for this model. This conclusion can also be made when comparing the values of the validation $\chi^{2}$ at fixed $\chi^{2}$. For $q=\pi/2$, the edge clearly preforms the best of the three parameterizations during the entire annealing run. For $q=3\pi/4$, the validation $\chi^{2}$ for the edge parametrization only dips below that of the other two at the end of the annealing run, but clearly reaches the lowest minimum value. We thus find that it is important take into consideration both the minimum value of each validation $\chi^{2}$ curve and the path the curves take when comparing the various parameterizations.
@@
We now compare the corresponding spectra, shown in Fig. 11, to see if the hierarchy suggested by the validation $\chi^{2}$ is borne out in the results of the analytic continuation process. As reflected by the relative differences between the validation $\chi^{2}$ curves, the single-peak and edge spectra share similar features, such as lower bounds near $\omega=1.5$ ($q=\pi/2$) and $\omega=1.0$ ($q=3\pi/4$), while the unconstrained sampling produces spectra with considerable spectral weight below the edge. The BA results [34], shown in black, agree very well with the edge spectra (with edges that nearly perfectly align), moderately well for the peak spectra (but with some large deviations), and very poorly with the unconstrained spectra, supporting the results of our cross validation test. Here, it should be noted that the BA spectra have been subject to broadening, and furthermore, the calculation of these spectra only take into consideration two-spinon and four-spinon contributions [35]. As such, it is known that a few percent of the total spectral weight is missing, which could at least partially explain the minor deviations between the two spectra in Fig. 11(c) and (f).
@@
The fact that the edge outperforms the unconstrained and single-peak parameterization strongly supports the viability of cross validation as a tool for model selection. Considering the fact that for a chain of length $512$, the exact $T=0$ spectrum likely only contains $\sim$100 delta functions with significant weight, deduced by extrapolating from the results for smaller system sizes [36, 37], it is remarkable that our method is able correctly identify the true form of the spectral function in the thermodynamic limit. It is quite possible that if smaller system size was considered instead, a spectrum composed of a dominate peak plus a smooth continuation would actually be the best representation and the single-peak parameterization would preform the best. But this is not the case here, implying that finite size effects are sufficiently mitigated.
VI.2 Heisenberg Chain with Long-Range Interactions